# Remove point-mass concept - remove singularities from GR

### v2

@@ Abstract
Keywords: Gravity beyond GR, Non-singular blackholes, Tully-Fisher relation, Einstein-Klein-Gordon equations, Flat rotation curves, Massive gravitons
PACS: $04.50.Kd$ Modified theories of gravity; $04.60.Bc$ Phenomenology of quantum gravity; $04.70.-s$ Physics of blackholes
I Introduction
In Newton’s gravitation, $GM/r$, or in Coulomb’s law, $Q/r$, one implicitly assumes there are physical entities of finite mass or charge packed in zero volumes. This assumption further implies that these point-likes can be assigned precise coordinates and momenta, and when approached infinitely closely, produce infinite gravitation or electric fields. That Newton’s and Coulomb’s laws have played decisive roles in setting the physics of the 17th-19th Centuries on axiomatized mathematical foundations is well acknowledged. That they are remarkably accurate to meet the everyday needs of the 21st century, from laboratory measurements to the solar system and galactic observations, is remarkable. The notion of point-like entities, however, is at odds with the foundations of quantum mechanics. Heisenberg’s uncertainty principle does not allow the assignment of precise coordinates and/or momenta to a physical entity. All one can hope for, is to find the particle, or whatever it is, in a quantum cloud of probability in a phase space.
@@ I Introduction
This paper builds upon our ongoing effort to address the limitations of the point-particle concept. In (Sobouti:2023kkb, ), one of us removes the Coulomb singularity of the Dirac electron by proposing a mutual action-reaction partnership between the Dirac wave function and the electric field of the electron itself. By so doing he comes up with a distributed charge and current for the spinning electron and the correct gyromagnetic ratio, however, without recourse to the QED formalism. In the case of GR, we argue that the classical energy-momentum tensor in Einstein’s field equation, no matter how one conceives it, has the point particle singularity inherent to it. In compliance with the uncertainty principle, we suggest to replace the classical energy-momentum tensor in Einstein’s field equation by a quantum mechanical equivalent, a collection of bosons each represented by a KG or by a normalizable YKG field.
@@
Resorting to auxiliary fields in GR, for different needs and purposes, has a long history and rich literature. Pioneering work by Ruffini and Bonazzola Ruffini:1969qy investigated systems of self-gravitating scalar bosons and spin-$1/2$ fermions. Based on numerical solutions, they conclude that spacetime remains non-singular. Buchdal Buchdahl:1959zz and Virbhadra Virbhadra:1997ie study the coupled Einstein-KG equations. Moffat Moffat:2005si and Moffat and Toth Moffat:2014pia propose TeVeS fields to have a modified gravity. Jetzer et al., Jetzer:1992tog , Bezares et al., Bezares:2022obu , Liddle and Madsen Liddle:1992fmk introduce scalar fields to have boson stars.
@@
What makes this paper different from the ones quoted above or not, is our emphasis on the fact that individual particles shaping the spacetime structure are extended wave-particle packets. We find that inclusion of this wave nature not only removes the essential and coordinate singularities, but alters the spacetime structure at galactic distances. Evidently the cumulative quantum waves of a degenerate and/or random collection of gravitating particles do not cancel out each other. At far outreaches of a galaxy, say, they create $r^{-1}$ force and cause flat rotation curves. We will come back to this issue in sections III.6 and V.
II Definition of the problem
@@
A pair of neutrons coupled through their isospin behave as an electrically neutral boson of spin $1$ or 00. See e.g. (Georgi:2000vve, ), chapter 5, section 5.4, Isospin generators . To mimic a neutron star one may consider a collection of such neutron nuggets. Neutrons aside, in a much wider context, one knows that about $99\%$ of the content of the universe is hydrogen and helium. They are either spread out in interstellar and intergalactic spaces or are in the form of ionized particles in an electrically neutral matrix of plasma in the stars. In the absence of any better choice and with some condone we contend to consider the matter we are dealing with as a collection of $N$ bosons in which each constituent is represented, for the moment, by a KG wave. For the total and individual Lagrangian density of the constituent bosons we write,
${\cal L}$ $=$ $\sum_{i=1}^{N}{\cal L}(i).$ (1)
${\cal L}_{i}$ $=$ $\frac{1}{2}\left[\frac{\hbar^{2}}{m}g^{\mu\nu}\partial_{\mu}\psi^{*}(i)\partial_{\nu}\psi(i)-mc^{2}\psi^{*}(i)\psi(i)\right],$ (2)
where $m$ is the mass of a single constituent, of the order of one or two nucleon mass. The Euler-Lagrange equation associated with (2) is,
@@
$=$ $-\frac{\hbar^{2}}{m}\partial_{\mu}\psi^{*}\partial_{\nu}\psi+\frac{1}{2}g_{\mu\nu}\left[\frac{\hbar^{2}}{m}\partial^{\alpha}\psi^{*}\partial_{\alpha}\psi-mc^{2}\psi^{*}\psi\right],$
@@
Einstein’s field equation for the collection of $N$ boson may now be written as,
$R_{\mu\nu}$ $=$ $\frac{8\pi G}{c^{4}}N\left[T_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T\right],$ (7)
$=$ $\frac{4\pi GNm}{c^{2}}\left[g_{\mu\nu}\psi^{*}\psi-2\frac{\hbar^{2}}{m^{2}c^{2}}\partial_{\mu}\psi^{*}\partial_{\nu}\psi\right]$
@@
In (9) and (II), $(\hbar^{2}\omega^{2}/m^{2}c^{4})$ is the square of oscillation energy to the rest mass energy of a KG wave. And $(\frac{\hbar^{2}}{m^{2}c^{2}}\approx 10^{-31}\textrm{m}^{2})$ is the square of the Compton wavelength of the mass $m$. Both are extremely small. Unless one is dealing with extreme relativistic cases, one may safely ignore them without compromising the essential role of the wave nature of the KG field. Furthermore, it is preferable to work with dimensionless quantities. We adopt
@@
Applying these simplifications to (9) and (II) we arrive at.
@@
$\frac{1}{x^{2}}\left[\frac{d}{dx}\left(\frac{x}{A}\right)-1\right]=\frac{2\pi N}{C}\psi^{*}\psi\,,$ (13)
@@ III Solution by Iteration
III.1 Near origin solutions, $x\rightarrow 0$.
@@
III.2 Near horizon solution, $x\rightarrow 1$
@@
III.3 Far distance solutions, $\mathrm{x\rightarrow\infty}$
@@
Such Yukawa amelioration is a logical and perhaps the imperative way out of the dilemma, for it is inconceivable to imagine that the wave function of a single particle extends to infinity and alters the spacetime structure at the scale of Universe.
@@
III.4 Time to revise our primitive model
In Section II we assumed the gravitating body consists of a collection of $N$ bosons, each represented by a KG, or now by a YKG, field. All constituents were assumed to be in their ground states. This is perhaps a good approximation for the core of those galaxies that have a central blackhole whose degenerate constituents are arranged in an orderly manner in the lowest quantum states in a coherent and phase-tuned manner (i.e. the same $\alpha$). The much larger fraction of the gravitating body, however, remains non-degenerate. Its constituents reside in the spacious phase space of states with no obligation to tune themselves with neighbors. Therefore, let us rewrite the Lagrangian of (2) as the sum of two degenerate and non-degenerate components:
@@ III.4 Time to revise our primitive model
where we have divided the gravitating matter into a fraction $f$ of degenerate core in their ground state and the remaining non-degenerate fraction $(1-f)$ in various $\psi(i)$ states (see Davis_2019 for blackholes at the center of spiral galaxies and the references therein. Typical values of $f$ are $10^{-3}-10^{-4}$). Constituents in the non-degenerate fraction will necessarily be in different non-correlated states and for all practical purposes will have random phases, $\alpha_{i}$. In line with this division of $\cal L$ the metric coefficients $A$ and $B$ divide accordingly. In particular the gravitational force $\frac{1}{2}c^{2}B^{\prime}$ of (27), that we will need it below, gets divided as follows,
$\frac{1}{2}c^{2}B^{\prime}$ $=$ $(fN)c^{2}e^{-2\kappa x}\Big{[}\frac{\kappa}{x}\Big{(}1-\frac{3}{4}\cos 2(kx+\alpha)\Big{)}+\frac{1}{2x^{2}}\Big{(}1+\frac{3}{4}\frac{\kappa}{k}\sin 2(kx+\alpha)\Big{)}\Big{]}$ (29)
$+$ $c^{2}\sum_{i}^{(1-f)N}e^{-2\kappa|\mathbf{x}-\mathbf{x_{i}}|}\Big{[}\frac{\kappa}{|\mathbf{x}-\mathbf{x}_{i}|}\Big{(}1-\frac{3}{4}\cos 2(k|\mathbf{x}-\mathbf{x}_{i}|+\alpha_{i})\Big{)}$
$\hskip 142.26378pt+\frac{1}{2|\mathbf{x}-\mathbf{x_{i}}|^{2}}\Big{(}1+\frac{3}{4}\frac{\kappa}{k}\sin 2(k|\mathbf{x}-\mathbf{x}_{i}|+\alpha_{i})\Big{)}\Big{]},$
where $x$ and $x_{i}$ are coordinates of the observation point and the source ones, respectively. Both $x_{i}$ and $\alpha_{i}$ are random variables and $(1-f)N$ is a huge number. At far distances, ($\mathbf{x}\gg\mathbf{x}_{i},\forall\mathbf{x}_{i}$), one may Taylor expand terms of the form $f(|\mathbf{x}-\mathbf{x}_{i}|)$ as follows
@@
Summing over $i$ eliminates all terms odd in $x_{i}$ due to their random nature. The next even term is an order of magnitude $(|\mathbf{x}_{i}|/|\mathbf{x}|)^{2}$ smaller than the first term and can be dropped as $x\rightarrow 0$. With the two trigonometric terms in (29), however, one should be careful. They are high-frequency random undulations. The routine to deal with them is to square random terms, sum them up, and take the square root of the sum. Reducing (29) as explained gives,
$\frac{1}{2}c^{2}B^{\prime}$ $=$ $fNc^{2}e^{-2\kappa x}\bigg{(}\frac{\kappa}{x}\Big{(}1-\frac{3}{4}\cos 2(kx+\alpha)\Big{)}+\frac{1}{2x^{2}}\Big{(}1+\frac{3}{4}\frac{\kappa}{k}\sin 2(kx+\alpha)\Big{)}\bigg{)}$ (30)
$+(1-f)Nc^{2}e^{-2\kappa x}\bigg{(}\frac{\kappa}{x}+\frac{1}{2x^{2}}\bigg{)}+\sqrt{(1-f)N}e^{-2\kappa x}\bigg{(}\frac{3}{8}\frac{\kappa}{x}+\frac{3}{16}\frac{\kappa}{kx^{2}}\bigg{)}.$
To have a better physical picture let us go back and rewrite (30) in terms of $r=r_{s}x$ and $d/dr=r^{-1}_{s}d/dx,~{}r_{s}=2Gm/c^{2}$.
$\frac{1}{2}c^{2}\frac{dB}{dr}$ $=$ $(fN)c^{2}e^{-2\kappa r/r_{s}}\bigg{(}\frac{\kappa}{r}\Big{(}1-\frac{3}{4}\cos 2(kr/r_{s}+\alpha)\Big{)}+\frac{r_{s}}{2r^{2}}\Big{(}1+\frac{3}{4}\frac{\kappa}{k}\sin 2(kr/r_{s}+\alpha)\Big{)}\bigg{)}$ (31)
$+c^{2}{(1-f)N}e^{-2\kappa r/r_{s}}\bigg{(}\frac{\kappa}{r}+\frac{r_{s}}{2r^{2}}\bigg{)}+c^{2}\sqrt{(1-f)N}e^{-2\kappa r/r_{s}}\bigg{(}\frac{3}{8}\frac{\kappa}{r}+\frac{3}{16}\frac{\kappa r_{s}}{kr^{2}}\bigg{)}.$
In (31), the sum of three $r^{-2}$ terms (ignoring the exponential falloff factor and $\kappa,~{}k$ corrections) is $c^{2}Nr_{s}/2r^{2}=GM/r^{2}$. It is the familiar Newtonian gravity and plays its dominant role at close and intermediate distances. The sum of three $r^{-1}$ terms, $c^{2}N\kappa/r$, are non-Newtonian and non-GR forces. They die out much slower than the Newtonian ones. We will shortly see that they are responsible for the flat rotation curves of the spiral galaxies and can be interpreted as an alternative gravity, a dark matter scenarios, etc. Sinusoidal terms come from the degenerate component and are indications of existence of non-singular blackholes in the system.
III.5 Determination of $\kappa$ Tully- Fisher relation
Observed facts and implication: In Newtonian gravitation the speed of a test object in circular orbit about a localized gravitating mass is $v^{2}=\frac{1}{2}c^{2}r\frac{dB}{dr}=\frac{GM}{r}$. This, however, is not the case for stars and HI clouds orbiting their host galaxies at large distances. The Tully-Fisher relation (TFr), sifted from a forage of observational data, Tully:1977fu , states that:

### v3

@@ Abstract
Dedicated to International Year of Quantum Science and Technology 2025.
[…]
PACS Nos.: $04.50.Kd$ Modified theories of gravity; $04.60.Bc$ Phenomenology of quantum gravity; $04.70.-s$ Physics of blackholes
keywords:
Gravity beyond GR; Non-singular blackholes; Tully-Fisher relation; Einstein-Klein-Gordon equations; Flat rotation curves; Massive gravitons.
1 Introduction
In Newton’s gravitation, $GM/r$, or in Coulomb’s law, $Q/r$, one implicitly assumes there are physical entities of finite mass or charge packed in zero volumes. The assumption further implies that these point-likes can be assigned precise coordinates and momenta, and when approached infinitely closely, produce infinite gravitation or electric fields. That Newton’s and Coulomb’s laws have played decisive roles in setting the physics of the 17th-19th Centuries on axiomatized mathematical foundations is well acknowledged. That they are remarkably accurate to meet the everyday needs of the 21st century, from laboratory measurements to the solar system and galactic observations, is remarkable. The notion of point-like entities, however, is at odds with the foundations of quantum mechanics. Heisenberg’s uncertainty principle does not allow the assignment of precise coordinates and/or momenta to a physical entity. All one can hope for, is to find the particle, or whatever it is, in a quantum cloud of probability in a phase space.
@@ 1 Introduction
This paper builds upon our ongoing effort to address the limitations of the point-particle concept. In [1], one of us removes the Coulomb singularity of the Dirac electron by proposing a mutual action-reaction partnership between the Dirac wave function and the electric field of the electron itself. By so doing he comes up with a distributed charge and current for the spinning electron and the correct gyromagnetic ratio, however, without recourse to the QED formalism. In the case of GR, we argue that the classical energy-momentum tensor in Einstein’s field equation, no matter how one conceives it, is a collection of inherently singular point particles. In compliance with the uncertainty principle, we suggest to replace (not the lumpsum of the gravitating matter but) each member in that collection of singulars by a quantum mechanical equivalent each represented by a KG or by a normalizable YKG field, say.
@@
Resorting to auxiliary fields in GR, for different needs and purposes, has a long history and rich literature. Pioneering work by Ruffini and Bonazzola [2] investigated systems of self-gravitating scalar bosons and spin-$1/2$ fermions. Based on numerical solutions, they conclude that spacetime remains non-singular. Buchdal [3] and Virbhadra [4] study the coupled Einstein-KG equations. Moffat [5] and Moffat and Toth [6] propose TeVeS fields to have a modified gravity. Jetzer et al. [7], Bezares et al. [8], Liddle and Madsen [9] introduce scalar fields to have boson stars. There are also a large number of papers on regular blackholes introduced since late 1960s, [10, 11] and onwards, see [12] for a short review. We remind that what we do here has no kinship with regular blackholes, where to avoid singularities one resorts to auxiliary energy momentum tensors.
@@
What differentiates this paper from the ones cited above or not is our focus on the fact that particles shaping the structure of the spacetime are not point-like but extended wave packets. We find that due recognition of this wave nature of individual gravitating particles not only removes the essential and coordinate singularities, but alters the spacetime structure at galactic distances. Evidently the cumulative quantum waves of a coherent and/or random collection of gravitating particles do not cancel out each other. At far outreaches of a galaxy, say, they create $r^{-1}$ force and cause flat rotation curves. We will come back to this issue in subsection 3.6 and section 5.
2 Definition of the problem
@@
A pair of neutrons coupled through their isospin behave as an electrically neutral boson of spin $1$ or 00. See e.g. [13], chapter 5, section 5.4, Isospin generators . To mimic a neutron star one may consider a collection of such neutron nuggets. Neutrons aside, in a much wider context, one knows that about $99\%$ of the content of the universe is hydrogen and helium. They are either spread out in interstellar and intergalactic spaces or are in the form of ionized particles in an electrically neutral matrix of plasma in the stars. In the absence of any better choice and with some condone we contend to consider the matter we are dealing with as a collection of $N$ bosons in which each constituent is represented, for the moment, by a KG wave. For the total and individual Lagrangian density of the constituent bosons we write,
${{\cal L}}$ $=$ $\sum_{i=1}^{N}{\cal L}_{i}\,,$ (1)
${\cal L}_{i}$ $=$ $\frac{1}{2}\left(\frac{\hbar^{2}}{m_{i}}g^{\mu\nu}\partial_{\mu}\psi^{*}(i)\partial_{\nu}\psi(i)-m_{i}c^{2}\psi^{*}(i)\psi(i)\right)\,,$ (2)
$=$ ${-{\frac{1}{2}}\psi^{*}\Big{(}\frac{{{\hbar^{2}}}}{{{m_{i}}}}\big{(}{g^{\mu\nu}}{\partial_{\mu}}{\partial_{\nu}}\psi(i)\big{)}+{{m_{i}}{c^{2}}\psi(i)}\Big{)}}$ (3)
$+$ ${\frac{1}{2}\frac{{{\hbar^{2}}}}{{{m_{i}}}}{\nabla_{\mu}}\Big{(}{g^{\mu\nu}}{\psi^{*}}(i){\partial_{\nu}}\psi(i)\Big{)}}$ (4)
where $m_{i}$ is the mass of a single constituent, of the order of one or two nucleon mass. The expression in (4) is a total derivative and does not contribute to the equation of motion . The latter from either (2) or (3) is
@@
$=$ $-\frac{\hbar^{2}}{m}\partial_{\mu}\psi^{*}\partial_{\nu}\psi+\frac{1}{2}g_{\mu\nu}\left(\frac{\hbar^{2}}{m}\partial^{\alpha}\psi^{*}\partial_{\alpha}\psi-mc^{2}\psi^{*}\psi\right),$
@@
Note the conservation of the energy-momentum tensor for each gravitating KG wave. Einstein’s field equation for the collection of $N$ boson may now be written as,
$R_{\mu\nu}$ $=$ $\frac{8\pi G}{c^{4}}N\left(T_{\mu\nu}-\frac{1}{2}g_{\mu\nu}T\right),$ (9)
$=$ $\frac{4\pi GNm}{c^{2}}\left(g_{\mu\nu}\psi^{*}\psi-2\frac{\hbar^{2}}{m^{2}c^{2}}\partial_{\mu}\psi^{*}\partial_{\nu}\psi\right)$
@@
In (11) and (2), $(\hbar^{2}\omega^{2}/m^{2}c^{4})$ is the square of oscillation energy to the rest mass energy of a KG wave. And $(\frac{\hbar^{2}}{m^{2}c^{2}}\approx 10^{-31}\textrm{m}^{2})$ is the square of the Compton wavelength of the mass $m$. Both are extremely small. Unless one is dealing with extreme relativistic cases, one may safely ignore them without compromising the essential role of the wave nature of the KG field. Furthermore, it is preferable to work with dimensionless quantities. We adopt
@@
Applying these simplifications to (11) and (2) we arrive at.
@@
$\frac{1}{x^{2}}\left(\frac{d}{dx}\left(\frac{x}{A}\right)-1\right)=-\frac{2\pi N}{C}\psi^{*}\psi\,,$ (15)
@@ 3 Solution by Iteration
3.1 Near origin solutions, $x\rightarrow 0$.
@@
3.2 Near horizon solution, $x\rightarrow 1$
@@
At this stage it is appropriate to pay tribute to Roy Kerr. In a recent paper [14] he expresses reservations on the singularity theorems of Penrose [15] and of Hawking [16]: ”there are no proofs that black holes contain singularities when they are generated by real physical bodies”. In his concluding remarks Roy further reiterates: ”The author has no doubt that, and he never did, that when Relativity and Quantum Mechanics are melded it will be shown that there are no singularities anywhere.”
What we conclude above is in agreement with Roy Kerr’s prophecy. Evidently the sole replacement of point-like gravitating masses by wave like ones is sufficient to wipe out both intrinsic and coordinate singularities from GR.
3.3 Far distance solutions, $\mathrm{x\rightarrow\infty}$
@@
This equation can be derived from:
${{\cal L}_{i}(YKG)}$ $=$ ${-{\frac{1}{2}}\psi^{*}\Big{(}\big{(}{g^{\mu\nu}}{\partial_{\mu}}{\partial_{\nu}}\psi(i)\big{)}}-{(\kappa\pm ik)^{2}(i)\psi(i)\Big{)}}\,.$ (25)
The Lagrangian (25) is the same as that in (3), simplified to account for the flatness of space-time and generalized to accommodate the Yukawa falloff parameter $\kappa$. Such Yukawa amelioration is a logical and perhaps the imperative way out of the dilemma, for it is inconceivable to imagine that the wave function of a single particle extends to infinity and alters the spacetime structure at the scale of Universe.
@@
3.4 Time to revise our primitive model
In Section 2 we assumed the gravitating body consists of a collection of $N$ bosons, each represented by a KG, or now by a YKG, field. All constituents were assumed to be in their ground states. This is perhaps a good approximation for the core of those galaxies that have a central blackhole whose degenerate constituents are arranged in an orderly manner in the lowest quantum states in a coherent and phase-tuned manner (i.e. the same $\alpha$). The much larger fraction of the gravitating body, however, remains non-degenerate. Its constituents reside in the spacious phase space of states with no obligation to tune themselves with neighbors. Therefore, let us rewrite the Lagrangian of (2) as the sum of two degenerate and non-degenerate components:
@@ 3.4 Time to revise our primitive model
where we have divided the gravitating matter into a fraction $f$ of degenerate core in their ground state and the remaining non-degenerate fraction $(1-f)$ in various $\psi(i)$ states (see [17] for blackholes at the center of spiral galaxies and the references therein. Typical values of $f$ are $10^{-3}-10^{-4}$). Constituents in the non-degenerate fraction will necessarily be in different non-correlated states and for all practical purposes will have random phases, $\alpha_{i}$. In line with this division of $\cal L$ the metric coefficients $A$ and $B$ divide accordingly. For instance $A$ of (28) gets divided as follows
$A(x)$ $=$ $1+f\frac{N}{{2x}}{e^{-2\kappa x}}\Big{(}1-\frac{\kappa}{k}\sin 2(kx+\alpha)\Big{)}$ (32)
$+$ $\sum\limits_{i}^{(1-f)N}{\frac{1}{{\left|{\mathbf{x}-{\mathbf{x}_{i}}}\right|}}}{e^{-2\kappa\left|{\mathbf{x}-{\mathbf{x}_{i}}}\right|}}\Big{(}1-\frac{\kappa}{k}\sin 2(k\left|{\mathbf{x}-{\mathbf{x}_{i}}}\right|+{\alpha_{i}})\Big{)},$
where x and $\textbf{x}_{i}$ are coordinates of the observation point and the source ones, respectively. Both $\textbf{x}_{i}$ and $\alpha_{i}$ are random variables and $(1-f)N$ is a huge number. At far distances, ($|\mathbf{x}|\gg|\mathbf{x}_{i}|,\forall|\mathbf{x}_{i}|$), one may Taylor expand terms of the form $f(|\mathbf{x}-\mathbf{x}_{i}|)$ as follows
@@
Summing over $i$ eliminates all terms odd in $x_{i}$ due to their random nature. The next even term is an order of magnitude $(|\mathbf{x}_{i}|/|\mathbf{x}|)^{2}$ smaller than the first term and can be dropped as $x\rightarrow 0$ (monopole approximation). With the two trigonometric terms in (32), however, one should be careful. They are high-frequency random undulations. The routine to deal with them is to square random terms, sum them up, and take the square root of the sum. Reducing (32) as explained gives,
$A(x)$ $=$ $1+f\frac{N}{{2x}}{e^{-2\kappa x}}\Big{(}1-\frac{\kappa}{k}\sin 2(kx+\alpha)\Big{)}$ (33)
$+$ $(1-f)\frac{N}{{2x}}{e^{-2\kappa x}}-\frac{\kappa}{{4kx}}\sqrt{(1-f)N}{e^{-2\kappa x}}\,,$
as noted above the factor $(1-f)N$ is a huge number. One may drop its square root and arrive at
$A(x)=1+\frac{N}{{2x}}{e^{-2\kappa x}}\left(1-f\frac{\kappa}{k}\sin 2(kx+\alpha)\right)\,.$ (34)
For the $B$ of (29), one similarly finds
$B(x)$ $=$ $1+2\frac{{\Phi(x)}}{{{c^{2}}}}$ (35)
$=$ $1-\frac{N}{{2x}}{e^{-2\kappa x}}\left(1+\frac{{3\kappa}}{{4k}}f\sin 2(kx+\alpha)\right)\,,$
where in the weak field approximation,
$\Phi(x)$ $=$ $\Phi(r/r_{s})=\frac{1}{2}c^{2}(B-1)$ (36)
$=$ $-\frac{Nc^{2}}{{4x}}{e^{-2\kappa x}}\left(1+\frac{{3\kappa}}{{4k}}f\sin 2(kx+\alpha)\right),$
is effectively the gravitational potential. Accordingly the gravitational force field becomes
$g(r)$ $=$ $-\frac{1}{2}{c^{2}}\frac{{dB}}{{dr}}=-\frac{1}{2}\frac{{{c^{2}}}}{{{r_{s}}}}\frac{{dB}}{{dx}}$ (37)
$\approx$ $\frac{1}{2}\frac{{{c^{2}}}}{{{r_{s}}}}{e^{-2\kappa r/r_{s}}}\left(\frac{N}{{2{x^{2}}}}+\frac{{2N\kappa}}{x}\big{(}1+\frac{3}{2}f\cos 2(kx+\alpha)\big{)}\right),$
$\approx$ ${e^{-2\kappa r/r_{s}}}\left(\frac{{GNm}}{{{r^{2}}}}+\frac{{N{c^{2}}\kappa}}{r}\big{(}1+\frac{3}{2}f\cos 2(kr/r_{s}+\alpha)\big{)}\right).$
Ignoring the exponential factor in (37), the $\frac{1}{r^{2}}$ term is the familiar Newtonian gravitational force. The $\frac{1}{r}$ term is non-Newtonian. It is the collective quantum effect of $N$ bosons in the model ($N=10^{69}$ nucleons in a Milky Way type galaxy). It falls off at much slower rate than the Newtonian force and is essentially responsible for flat rotation curves in actual spiral galaxies. The sinusoidal factor in (37), proportional to $f$, comes from a would be central blackhole.
3.5 Determination of $\kappa$ Tully- Fisher relation
Observed facts and implication: In Newtonian gravitation the speed of a test object in circular orbit about a localized gravitating mass is $v^{2}=\frac{1}{2}c^{2}r\frac{dB}{dr}=\frac{GM}{r}$. This, however, is not the case for stars and HI clouds orbiting their host galaxies at large distances. The Tully-Fisher relation (TFr), sifted from a forage of observational data, [18], states that: